Transfer Pricing

Aulman Inc. has a number of divisions, including a Furniture Division and a Motel Division. The Motel Division owns and operates a line of budget motels located along major highways. Each year, the Motel Division purchases furniture for the motel rooms. Currently, it purchases a basic dresser from an outside supplier for $40. The manager of the Furniture Division has approached the manager of the Motel Division about selling dressers to the Motel Division. The full product cost of a dresser is $29. The Furniture Division can sell all of the dressers it makes to outside companies for $40. The Motel Division needs 10,000 dressers per year; the Furniture Division can make up to 50,000 dressers per year.

Also, although the Furniture Division has been operating at capacity (50,000 dressers per year), it expects to produce and sell only 40,000 dressers for $40 each next year. The Furniture Division incurs variable costs of $14 per dresser. The company policy is that all transfer prices are negotiated by the divisions involved.

Required:

1.  What is the maximum transfer price?
$

Which division sets it?
 

2.  What is the minimum transfer price?
$

Which division sets it?
 

3.  Suppose that the two divisions agree on a transfer price of $35. What is the benefit for the Furniture Division? For the Motel Division? For Aulman Inc. as a whole?

Benefit to Furniture Division$
Benefit to Motel Division$
Benefit to company$
Answer #1
Concepts and reason

Transfer Price: Transfer price is value for the inter-company or inter-departmental transfer of products. One department charges a transfer price for transferring its product to the other department. Transfer price is recorded for the accounting purpose. The performance measurement is done on the basis of this rate.

The performance evaluation and incentives of the division manager depend on the transfer value or the sales value of the products manufactured in their division. If products manufactured are transferred at the lower price, then the performance that division would be underestimated or lowered. This is the reason behind the disagreement of the managers in the same company over the transfer price.

Fundamentals

If the division has excess capacity available, then it transfers the product at the minimum price, which is the variable cost of production.

If the division does not have excess capacity available, then it transfers the product at the market price or variable cost of production per unit plus opportunity cost.

(a)

Determine the maximum transfer price as follows: -

The maximum transfer price is $40 per unit.

Therefore, the maximum transfer price is $40 per unit.

(b)

Determine the minimum transfer price per unit as follows: -

The minimum transfer price per unit is $14.

Therefore, the minimum transfer price per unit is $14.

(c)

Transfer price per unit is $35 from division F to division M. Total transferred unit is 10,000.

Benefit to Division F, M, and company is $210,000; $50,000 and $260,000, respectively.

Working note:

Determine the benefit to division F as follows: -

Benefit=(TransferPriceVariablecost)×TransferredUnits=($35$14)×10,000=$210,000\begin{array}{c}\\Benefit\, = \,(Transfer\,\Pr ice\, - Variable\,\,\cos t)\, \times Transferred\,Units\\\\ = \,(\$ 35 - \$ 14)\, \times 10,000\\\\ = \,\$ 210,000\\\end{array}

Therefore, the benefit to division F is $210,000.

Determine the benefit to division M as follows: -

Benefit=(MarketPriceTransferPrice)×TransferredUnits=($40$35)×10,000=$50,000\begin{array}{c}\\Benefit\, = \,(Market\,\,\Pr ice\, - Transfer\,\Pr ice)\, \times Transferred\,Units\\\\ = \,(\$ 40 - \$ 35)\, \times 10,000\\\\ = \,\$ 50,000\\\end{array}

Therefore, the benefit to division M is $50,000.

Determine the benefit to the company as follows: -

Benefit=BenefittoF+BenefittoM=$210,000+$50,000=$260,000\begin{array}{c}\\Benefit\, = \,Benefit\,to\,F + Benefit\,to\,M\\\\ = \,\$ 210,000 + \$ 50,000\\\\ = \,\$ 260,000\\\end{array}

Therefore, the benefit to the company is $260,000.

Ans: Part a

The maximum transfer price is $40 per unit.

Part b

The minimum transfer price per unit is $14.

Part c

Benefit to Division F, M, and company is $210,000; $50,000 and $260,000, respectively.